OpenStudy (anonymous):

write each repeating decimal as a fraction in simplest form. (a)0.16666666666666666666666666666666......66666666..... (b)o.4166666666666666666666666666666.....666666666666.....

OpenStudy (anonymous):

let's do the first one and you can do the same with the other...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

let x = 0.16666... now, 10x = 1.6666.... and 100x = 16.6666.. notice that 100x - 10x = 90x and 16.6666.... - 1.6666... = 15 so we have 90x = 15 or x=15/90... reduce that fraction....

OpenStudy (anonymous):

do you need anything clarified?

OpenStudy (anonymous):

um, I kind of get thegist of it, but I don't understand... you don't need to do the infinite sum of a geometric series?

OpenStudy (anonymous):

This one of those things that you can do mentally:\[ 0.4166666666 = \frac{416-41}{900} = \frac 5 {12} \]

OpenStudy (anonymous):

so what if it asks for u to solve it wth the infiinite sum of a geometric series way?

OpenStudy (anonymous):

That's how @dpaInc solved it.

OpenStudy (anonymous):

so this method IS the sum of an Infinite series way?

OpenStudy (anonymous):

0.1 + 6/100 + 6/1000 + 6/10000 +... so 1/10 + sum of infinite geom. series where an = 6*(1/10)^(n+1)) a = 6/100 and r = 1/10 sum = a/(1-r) = 6/100 / (1 - 1/10) = 6/100 / 9/10 = 6/100 * 10/9 =6/90 NOT done... add this to 1/10 1/10 + 6/90 = 15/90 now reduce.... Here is a link: http://en.wikipedia.org/wiki/Geometric_series More than one way to do this.

OpenStudy (anonymous):

thank you johnrr!